Show that the points (2,0,1),(0,-4,3) and (-2,5,0) are non-collinear. Hence, find the equation of plane passing through them
1
Expert's answer
2020-08-10T16:17:04-0400

Let check the determinant is non zero or not


"\begin{vmatrix}
 2 & 0&1 \\
 0 & -4&3\\
-2&5&0
\end{vmatrix}=-19\neq0"

Thus points are non collinear.

Equation of plane:

"\begin{vmatrix}
 x-2&y-0&z-1\\
2-0&0-(-4)&1-3
\\
0-(-2)&-4-5&3-0
\end{vmatrix}=\begin{vmatrix}
 x-2&y-0&z-1\\
2&4&-2
\\
2&-9&3
\end{vmatrix}=0\\
\implies-6(x-2)-10(y-0)+-26(z-1)=0\\\implies -6x-10y-26z+38=0"